OpenStudy (anonymous):

Solve by factoring 9x^2+30x+25?

OpenStudy (gw2011):

(3x+5)(3x+5) If you are trying to solve for x, then you need to set the factors to zero: (3x+5)(3x+5)=0 3x+5=0 3x=-5 x=-5/3

OpenStudy (anonymous):

How did you get (3x+5)?

OpenStudy (gw2011):

You need to find what will factor. By looking at the problem you can see that (3x)(3x)=9x^2 and that (5)(5)=25. Using these you to need to make sure that the middle term will add to 30x. Therefore, (3x)(5)+(3x)(5)=15x+15x=30x. As a result the factor comes out to be: (3x+5)(3x+5).